An online Merchant estimates the price elasticity for rice is estimated to be -0.4 and the income elasticity is 0.8. At a price of $0.40 per pound and a per capita income of $20,000, the demand for the rice is 50million tons per year.
Required;
a) If per capita increases to $20,500, what will be the quantity demanded of rice.
b) If price of rice increases to $0.41 per pound and income remains at $20,000, what will be the quantity demanded.
1
Expert's answer
2019-11-12T17:00:24-0500

a) If per capita increases to $20,500, then the quantity demanded of rice will increase by (20,500/20,000 - 1)×0.8 = 0.02 or 2% to 50×1.02 = 51 million tons.

b) If price of rice increases to $0.41 per pound and income remains at $20,000, then the the quantity demanded will decrease by (0.41/0.40 - 1)×(-0.4) = -0.01 or -1% to 50×0.99 = 49.5 million tons.